# Dual chiral density wave induced oscillating Casimir effect

### v1

@@ I Introduction
Beyond academic interest, the engineering application of the Casimir effect to nanotechnology (Casimir engineering) has recently attracted much attention [9]. The most typical feature of the Casimir effect is an attractive force. On the other hand, when one tunes the permittivity and/or permeability of plates and medium, a repulsive force can be also realized [10, 11, 12, 13]. In contrast to such attractive or repulsive Casimir effects, it would be interesting that the third type of Casimir effect, where “third” means that attraction or repulsion is not fixed. For example, under a setup, the sign of Casimir energy flips from attraction to repulsion as the separation distance increases, which may be called the sign-flipping Casimir effect (e.g., see Refs. [14, 15, 16, 17, 18, 19]). As another example, the value of Casimir energy can oscillate as a function of distance, which may be called the oscillating Casimir effect (e.g., see Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]). Among them, Ref. [31] found that it occurs inside Weyl semimetals, where the origins of the oscillation are Weyl points (WPs) at finite momenta in the dispersion relations of Weyl fermions. Such new types of Casimir effects are not only of theoretical interest but also will be important for Casimir engineering.
@@
Under these motivations, this paper focuses on what types of Casimir effects can occur in various phases of quark matter where the thickness of $z$ direction is extremely short (i.e., “thin” quark matter) as illustrated in Fig. 1. In particular, in this paper, we propose a QCD counterpart of the oscillating Casimir effect predicted in Weyl semimetals. This is realized for fermion fields in the dual chiral density wave (DCDW) phase [33] (also see Refs. [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] and Ref. [81] for a review). The DCDW phase is a candidate for the ground state of quark/nuclear matter in the density region of $\rho/\rho_{0}=3\sim 6$ with the normal nuclear density $\rho_{0}=0.16$ ${\rm fm}^{-3}$ [46, 47], and complex dispersion relations of fermions in this phase can be an origin of the oscillating Casimir effect.
@@ II.3 Casimir energy with Lifshitz formula
The Casimir energy is defined as a finite-volume effect for the zero-point energy (10). In a finite volume, the momentum integral with respect to the three-dimensional momentum of quarks in Eq. (10) is replaced by a discrete sum. In this paper, we impose the periodic boundary conditions (PBC) on quark fields at $z=0$ and $z=L_{z}$, where the $z$ component of quark momentum is discretized as $k_{z}\rightarrow 2n\pi/L_{z}$ where $n=0,1,\cdots,\infty$. [^2] The case with boundary conditions for the $x$ or $y$ direction is straightforward, but these boundaries do not induce the WP-induced oscillating Casimir effect since now the DCDW is along the $z$ axis. Note that for the PBC, there is no outside vacuum as in Fig. 1, but the discretization of momentum (and corresponding eigenvalues) can induce a nonzero Casimir energy.
@@
$E_{\rm Cas}$ $=-4N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{dk_{x}dk_{y}}{(2\pi)^{2}}\ln\big{[}1-e^{-L_{z}k_{z}^{[s]}}\big{]},$
$k^{[\pm]}_{z}$ $=\sqrt{k_{\perp}^{2}+M^{2}+\xi^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k_{\perp}^{2}+\xi^{2}}},$ (11)
where the first integral variable is the imaginary part $\xi$ of imaginary frequency $\omega\equiv i\xi$, and $k_{\perp}^{2}\equiv k_{x}^{2}+k_{y}^{2}$. The overall factor of $2N_{f}N_{c}$ means the degrees of freedom for particle-antiparticle, flavors, and colors, and the spin degrees of freedom are labeled as $s=\pm$. The overall factor of $-2$ and the factor of $-e^{-L_{z}k_{z}^{[s]}}$ are a property of the PBC on fermion fields. For the MIT bag boundary condition, see Appendix A.
Note that Eq. (11) is one of the main findings in this paper, which is an analog of the Lifshitz formula [88] to calculate the Casimir effect for photon fields. When we substitute $M=q=0$ and $q=0$, Eq. (11) returns to the formulas for the usual massless and massive quarks, respectively.
@@
Figure 2: (a) Dispersion relations of the four eigenmodes, Eqs. (6) and (7), at $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.1,0.6,0,6)$. (b) Each contribution from the Dirac sea (red line) or the Fermi sea (blue line) and their sum (purple line).
@@ II.4 Casimir energy with a lattice regularization
where the dispersion relations $E^{\rm Lat}_{\pm}$ on the lattice is obtained by replacing the original momenta $k^{2}_{i}$ ($i=x,y,z$) in Eqs. (6) and (7) as
@@
The first term $E^{\rm sum}_{0}$ in Eq. (13) is the zero-point energy in finite volume (and at finite chemical potential) with the momentum discretized by a finite thickness $L_{z}=aN_{z}$, where $N_{z}$ is the number of lattice cells. The second term $E^{\rm int}_{0}$ in Eq. (13) is the zero-point energy with the continuous momentum in infinite volume. The Casimir energy $E^{\rm Lat}_{\rm Cas}$ is defined as the difference between these two energies. Also, as a result of lattice regularization, the momentum integral and the momentum sum are taken within the first Brillouin zone (BZ). In the case of the PBC, the sum is over $n=0,1,...,N_{z}-1$ (or equivalently $n=0,1,...,N_{z}$).
@@ II.5 Dispersion relations
Figure 2(b) shows each contribution from the Dirac or Fermi sea and their sum. The Dirac-sea and Fermi-sea contributions correspond to the absolute value of the first and second terms of the integrand in Eq. (10), respectively. These dispersion relations become flat bands at low momentum and bend twice in the middle of the dispersion relation. Such non-differentiable points in the dispersion relation generally lead to an oscillating Casimir effect.
@@ III Results
In this section, we show the results for the Casimir coefficients in the low-, intermediate-, and high-density regions, with the PBC (see Appendix A for the discussion with the MIT bag boundary condition). In the following, with $\Lambda=860\ {\rm MeV}$, we fix the values of the order parameters, $M$ and $q$, at three $\mu$ (low-, intermediate, and high-density regions) as those obtained in Ref. [46], where the authors applied the mean-field approach in the NJL model with the proper-time regularization. [^3] Note that the values of order parameters, $M$ and $q$, in the NJL model depend on regularization schemes. The Casimir energy is a quantity independent of regularization schemes. Therefore, once $M$ and $q$ are obtained by a regularization, we can calculate the Casimir energy equivalently by any regularization (using the fixed $M$ and $q$).
@@ III.1 Oscillating Casimir effect (virtual parameters)
For example, in the low-momentum region of a massive particle, since the dispersion relation is a downward convex function, then $S^{\rm int}_{n}>S^{\rm sum}_{n}$ is satisfied, leading to a positive contribution to $N_{z}(S^{\rm sum}_{n}-S^{\rm int}_{n})\propto E^{\rm Lat}_{\rm Cas}$. [^4] This is a simplified situation. As a more complex case, when a dispersion relation is inverted by a Wely point or a Fermi point, the function in the high-momentum side can be an upward convex function. On the other hand, in the high-momentum region, since the dispersion relation is an upward convex function due to the lattice regularization, $S^{\rm sum}_{n}>S^{\rm int}_{n}$, which leads to a negative contribution to $E^{\rm Lat}_{\rm Cas}$. In a region that can be regarded as a linear dispersion, we have $S^{\rm sum}_{n}=S^{\rm int}_{n}$.
@@ III.2 Low-density region
Figure 5 shows the dispersion relation of the quark field along the momentum $k_{z}$ on the lattice with $a\Lambda=1$, where the blue line represents the contribution from the Dirac sea, the red line from the Fermi sea, and the purple line their sum.
Figure 5: Dispersion relations on the lattice with $a\Lambda=1$ in the low-density region ($\mu/\Lambda=0.46$).
In this case, there is no mode crossing the Fermi level, which means that the total dispersion relation has no non-differentiable point. Therefore, the four dispersion relations are nothing but that of a massive Dirac fermion.
@@
Figure 6: Casimir coefficients in the low-density region. We denote $a\Lambda$ as $a$ in the legends.
@@ III.3 Intermediate-density region
As the chemical potential increases, the contribution from the quark Fermi sea causes a significant change in the behavior of the Casimir effect. Figure 7 shows the dispersion relation of the quark field along the momentum $k_{z}$ for fixed parameters $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.46,0,0.485)$ and $a\Lambda=1$. In this parameter set, the DCDW phase has not yet occurred, but the $\omega_{\pm}$ with low momentum is below the Fermi sea. Then, there is an additional contribution to the Casimir effect from the Fermi sea as well as from the Dirac sea. Furthermore, the crossing points (Fermi points) of $\omega_{\pm}$ and the Fermi level lead to non-differentiable points in the total dispersion relation. Due to this, the oscillating Casimir effect occurs, by a mechanism similar to that discussed in the subsection II.5.
@@
Figure 8 shows the results for the Casimir coefficient. In the short thickness of Fig. 8, we find that the Casimir coefficient is dominated by the contribution from the Dirac sea, which is qualitatively the same as the result shown in the previous subsection III.2. On the other hand, in the long thickness of Fig. 8, the Casimir coefficient is dominated by the contribution from the Fermi sea. As a result, the sign of the total Casimir coefficient flips around $aN_{z}=L_{z}\sim 2.3\ {\rm fm}$ and $8.8\ {\rm fm}$. At further long-thickness, we expect an oscillation of the Casimir coefficient. However, now we do not show the numerical results due to the limitations of calculations with sufficient accuracy. By using Eq. (19), the period of oscillation is estimated to be $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 9.4\ {\rm fm}$, where $k_{\rm FP}/\Lambda\sim 0.154$ is the momentum of the Fermi point.
@@
When the chemical potential is large enough, the DCDW phase ($q\neq 0$) is realized. Figure 9 shows the dispersion relation of the quark fields along the momentum $k_{z}$ for fixed parameters $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.09,0.62,0.72)$ with $a\Lambda=1$.
Figure 9: Dispersion relations on the lattice with $a\Lambda=1$ in the high-density region ($\mu/\Lambda=0.72$).
In this case, the two touching points of $\omega_{-}$ and $\tilde{\omega}_{-}$ are regarded as the Weyl points. Furthermore, the $\omega_{+}$ and $\omega_{-}$ cross the Fermi level. Thus, there are six non-differentiable points. In the low-momentum region, the total dispersion relation becomes a flat band due to the cancellations between $\omega_{-}$ and $\tilde{\omega}_{-}$ and between $\omega_{+}$ and $\tilde{\omega}_{+}$, and the edges of the flat band of the total dispersion relation are non-differentiable. Furthermore, due to the four Fermi points, the total dispersion relation is non-differentiable at the four momenta.
Figure 10 shows the results for the Casimir coefficient.
@@
For the Dirac-sea effect, the Casimir effect shows oscillating behavior, as expected from the Weyl points. On the other hand, for the Fermi-sea effect as depicted by the red line, we see that the Fermi-sea part of Casimir coefficient oscillates with higher frequency than that from the Dirac sea. Its frequency is determined by the positions of the non-differentiable points on the dispersion relation. By using Eq. (19), the oscillation periods are estimated to be $L_{z}^{\rm osc}=2\pi/k_{\rm WP}\sim 4.9$ fm with $k_{\rm WP}/\Lambda\sim 0.297$ (for the Dirac sea part) and $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 1.4$ fm with $k_{\rm FP}/\Lambda\sim 1.03$ (for the Fermi sea part of $\omega_{-}$), and $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 3.6$ fm with $k_{\rm FP}/\Lambda\sim 0.40$ (for the Fermi sea part of $\omega_{+}$). The Dirac-sea and Fermi-sea contributions shown in Fig. 10 are dominated by the periods of $L_{z}^{\rm osc}\sim 4.9$ fm and $L_{z}^{\rm osc}\sim 1.4$ fm, respectively, and the total result is their superposition.

### v2

@@ I Introduction
Beyond academic interest, the engineering application of the Casimir effect to nanotechnology (Casimir engineering) has recently attracted much attention [9]. The most typical feature of the Casimir effect is an attractive force. On the other hand, when one tunes the permittivity and/or permeability of plates and medium, a repulsive force can be also realized [10, 11, 12, 13]. In contrast to such attractive or repulsive Casimir effects, the third type of Casimir effect would be interesting, where “third” means that attraction or repulsion is not fixed. For example, under a setup, the sign of Casimir energy flips from attraction to repulsion as the separation distance increases, which may be called the sign-flipping Casimir effect (e.g., see Refs. [14, 15, 16, 17, 18, 19]). As another example, the value of Casimir energy can oscillate as a function of distance, which may be called the oscillating Casimir effect (e.g., see Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]). Among them, Ref. [31] found that it occurs inside Weyl semimetals, where the origins of the oscillation are Weyl points (WPs) at finite momenta in the dispersion relations of Weyl fermions. Such new types of Casimir effects are not only of theoretical interest but also will be important for Casimir engineering.
@@
Under these motivations, this paper focuses on what types of Casimir effects can occur in various phases of quark matter where the thickness of $z$ direction is extremely short (i.e., “thin” quark matter) as illustrated in Fig. 1. In particular, in this paper, we propose a QCD counterpart of the oscillating Casimir effect predicted in Weyl semimetals. This is realized for fermion fields in the dual chiral density wave (DCDW) phase [33] (also see Refs. [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] and for related studies and Ref. [81] for a review). The DCDW phase is a candidate for the ground state of quark/nuclear matter in the density region of $\rho/\rho_{0}=3\sim 6$ with the normal nuclear density $\rho_{0}=0.16$ ${\rm fm}^{-3}$ [46, 47], and complex dispersion relations of fermions in this phase can be an origin of the oscillating Casimir effect.
@@ II.3 Casimir energy with Lifshitz formula
The Casimir energy is defined as a finite-volume effect for the zero-point energy (10). In a finite volume, the momentum integral with respect to the three-dimensional momentum of quarks in Eq. (10) is replaced by a discrete sum. In this paper, we impose the periodic boundary conditions (PBCs) on quark fields at $z=0$ and $z=L_{z}$, where the $z$ component of quark momentum is discretized as $k_{z}\rightarrow 2n\pi/L_{z}$ where $n=0,1,\cdots,\infty$. [^2] The case with boundary conditions for the $x$ or $y$ direction is straightforward, but these boundaries do not induce the WP-induced oscillating Casimir effect since now the DCDW is along the $z$ axis. Note that for the PBC, there is no outside vacuum as in Fig. 1, but the discretization of momentum (and corresponding eigenvalues) can induce a nonzero Casimir energy.
@@
$E_{\rm Cas}$ $=-4N_{f}N_{c}\sum_{s=\pm}\int_{0}^{\infty}\frac{d\xi}{2\pi}\int\frac{dk_{x}dk_{y}}{(2\pi)^{2}}\ln\big{[}1-e^{-L_{z}\tilde{k}_{z}^{[s]}}\big{]},$
$\tilde{k}^{[\pm]}_{z}$ $=\sqrt{k_{\perp}^{2}+M^{2}+\xi^{2}-\frac{q^{2}}{4}\mp iq\sqrt{k_{\perp}^{2}+\xi^{2}}},$ (11)
where the first integral variable is the imaginary part $\xi$ of imaginary frequency $\omega\equiv i\xi$, and $k_{\perp}^{2}\equiv k_{x}^{2}+k_{y}^{2}$. The overall factor of $2N_{f}N_{c}$ means the degrees of freedom for particle-antiparticle, flavors, and colors, and the spin degrees of freedom are labeled as $s=\pm$. The overall factor of $-2$ and the factor of $-e^{-L_{z}k_{z}^{[s]}}$ are a property of the PBC on fermion fields. See Appendix A for a derivation of Eq. (11) and the case of the antiperiodic boundary condition. Also, for discussion with the MIT bag boundary condition, see Appendix B.
Note that Eq. (11) is one of the main findings in this paper, which is an analog of the Lifshitz formula [88] to calculate the Casimir effect for photon fields. When we substitute $M=q=0$ and $q=0$, Eq. (11) returns to the known formulas for the massless and massive quarks, $E_{\rm Cas}=N_{f}N_{c}\times 2\pi^{2}/45L_{z}^{3}$ and $E_{\rm Cas}=N_{f}N_{c}\times(2M^{2}/\pi^{2}L_{z})\sum_{l=1}^{\infty}K_{2}(lML_{z})/l^{2}$, respectively, where $K_{2}$ is the modified Bessel function.
@@
Figure 2: (a) Dispersion relations of the four eigenmodes, Eqs. (6) and (7), at $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.1,0.6,0,6)$. The green region represents the Dirac and Fermi seas. (b) Each contribution from the Dirac sea (red line) or the Fermi sea (blue line) and their sum (purple line).
@@ II.4 Casimir energy with a lattice regularization
where the dispersion relations $E^{\rm Lat}_{\pm}$ on the lattice is obtained by replacing the original momenta $k^{2}_{i}$ ($i=x,y,z$) in Eq. (8) as
@@
The first term $E^{\rm sum}_{0}$ in Eq. (13) is the zero-point energy in finite volume (and at finite chemical potential) with the momentum discretized by a finite thickness $L_{z}=aN_{z}$, where $N_{z}$ is the number of lattice cells. The second term $E^{\rm int}_{0}$ in Eq. (13) is the zero-point energy with the continuous momentum in infinite volume. The Casimir energy $E^{\rm Lat}_{\rm Cas}$ is defined as the difference between these two energies. Also, as a result of lattice regularization, the momentum integral and the momentum sum are taken within the first Brillouin zone (BZ). In the case of the PBC, the sum is over $n=0,1,...,N_{z}-1$ (or equivalently $n=1,2,...,N_{z}$).
@@ II.5 Dispersion relations
Figure 2(b) shows each contribution from the Dirac or Fermi sea and their sum. The Dirac-sea and Fermi-sea contributions correspond to the absolute value of the first and second terms of the integrand in Eq. (10), respectively. These dispersion relations become flat bands at low momentum and bend twice in the middle of the dispersion relation. Such nondifferentiable points in the dispersion relation generally lead to an oscillating Casimir effect.
@@ III Results
In this section, we show the results for the Casimir coefficients in the zero-, intermediate-, and high-density regions, with the PBC (see Appendix B for the discussion with the MIT bag boundary condition). In the following, with $\Lambda=860\ {\rm MeV}$ and $1=\hbar c\sim 197.327\ {\rm MeV\cdot fm}$ ($\hbar$ is the reduced Planck constant, and $c$ the speed of light), we fix the values of the order parameters, $M$ and $q$, at three $\mu$ (zero-, intermediate-, and high-density regions) as those obtained in Ref. [46], where the authors applied the mean-field approach in the NJL model with the proper-time regularization. [^3] Note that the values of order parameters, $M$ and $q$, in the NJL model depend on regularization schemes. The Casimir energy is a quantity independent of regularization schemes. Therefore, once $M$ and $q$ are obtained by a regularization, we can calculate the Casimir energy equivalently by any regularization (using the fixed $M$ and $q$).
@@ III.1 Oscillating Casimir effect (virtual parameters)
For example, in the low-momentum region of a massive particle, since the dispersion relation is a downward convex function, then $S^{\rm int}_{n}>S^{\rm sum}_{n}$ is satisfied, leading to a positive contribution to $N_{z}(S^{\rm sum}_{n}-S^{\rm int}_{n})\propto E^{\rm Lat}_{\rm Cas}$. [^4] This is a simplified situation. As a more complex case, when a dispersion relation is inverted by a Weyl point or a Fermi point, the function in the high-momentum side can be an upward convex function. On the other hand, in the high-momentum region, since the dispersion relation is an upward convex function due to the lattice regularization, $S^{\rm sum}_{n}>S^{\rm int}_{n}$, which leads to a negative contribution to $E^{\rm Lat}_{\rm Cas}$. In a region that can be regarded as a linear dispersion, we have $S^{\rm sum}_{n}=S^{\rm int}_{n}$.
@@ III.2 Zero-density region
Figure 5 shows the dispersion relation of the quark field along the momentum $k_{z}$ on the lattice with $a\Lambda=1$, where the red line represents the contribution from the Dirac sea, the blue line from the Fermi sea, and the purple line from their sum.
Figure 5: Dispersion relations on the lattice with $a\Lambda=1$ in the zero-density region ($\mu/\Lambda=0.46$).
In this case, there is no mode crossing the Fermi level, which means that the total dispersion relation has no nondifferentiable point. Therefore, the four dispersion relations are nothing but that of a massive Dirac fermion.
@@
Figure 6: Casimir coefficients in the zero-density region. We denote $a\Lambda$ as $a$ in the legends.
@@
Finally, we comment on the physical scale of the vertical axis in Fig. 6. Since now we define the Casimir energy $E_{\mathrm{Cas}}$ as an energy per unit area, at $L_{z}=1$ fm in Fig. 6, $E_{\mathrm{Cas}}=\hbar c\times C_{\mathrm{Cas}}^{[3]}/L_{z}^{3}\sim 241$ $\mathrm{MeV}/\mathrm{fm}^{2}\sim 3.86\times 10^{4}$ $\mathrm{N}/\mathrm{fm}$. As a reference, for the case of massless quarks $(M=q=0)$, $C_{\mathrm{Cas}}^{[3]}=N_{f}N_{c}\times 2\pi^{2}/45\sim 2.63$ at any $L_{z}$. [^5] This $C_{\mathrm{Cas}}^{[3]}$ is equal to the value at $L_{z}=0$ in Fig. 6. This is because the effect of $M$ is small enough in $L_{z}\to 0$ due to $k_{z}^{2}\propto 1/L_{z}^{2}\gg M^{2}$. Then, at $L_{z}=1$ fm, $E_{\mathrm{Cas}}\sim 519$ $\mathrm{MeV}/\mathrm{fm}^{2}\sim 8.32\times 10^{4}$ $\mathrm{N}/\mathrm{fm}$ [^6] This is $-192$ times larger than the well-known value $E_{\mathrm{Cas}}=-\hbar c\frac{\pi^{2}}{720L_{z}^{3}}$ of the Casimir energy for the photon fields between perfectly conducting parallel plates. The factor of $-192$ comes from $2N_{f}N_{c}$ for the additional degrees of freedom and $16$ for the periodic boundary conditions. Note that the other interesting quantities are the Casimir pressure $P_{\mathrm{Cas}}$ and the Casimir force $F_{\mathrm{Cas}}$, which are defined as $P_{\mathrm{Cas}}\equiv F_{\mathrm{Cas}}/L_{x}L_{y}\equiv-\frac{d}{dL_{z}}E_{\mathrm{Cas}}$. As a reference, for the massless quarks at $L_{z}=1$ fm, $P_{\mathrm{Cas}}\sim 2.50\times 10^{5}$ $\mathrm{N}/\mathrm{fm}^{2}$.
@@ III.3 Intermediate-density region
As the chemical potential increases, the contribution from the quark Fermi sea causes a significant change in the behavior of the Casimir effect. Figure 7 shows the dispersion relation of the quark field along the momentum $k_{z}$ for fixed parameters $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.46,0,0.485)$ and $a\Lambda=1$. In this parameter set, the DCDW phase has not yet occurred, but the $\omega_{\pm}$ with low momentum is below the Fermi sea. Then, there is an additional contribution to the Casimir effect from the Fermi sea as well as from the Dirac sea. Furthermore, the crossing points (Fermi points) of $\omega_{\pm}$ and the Fermi level lead to nondifferentiable points in the total dispersion relation. Due to this, the oscillating Casimir effect occurs, by a mechanism similar to that discussed in Sec. III.1.
Figure 8 shows the results for the Casimir coefficient. In the short-thickness region, we find that the Casimir coefficient is dominated by the contribution from the Dirac sea, which is qualitatively the same as the result shown in the previous Sec. III.2. On the other hand, in the long-thickness region, the Casimir coefficient is dominated by the contribution from the Fermi sea. As a result, the sign of the total Casimir coefficient flips around $aN_{z}=L_{z}\sim 2.3\ {\rm fm}$ and $8.8\ {\rm fm}$. At further long thickness, we expect an oscillation of the Casimir coefficient. However, now we do not show the numerical results due to the limitations of calculations with sufficient accuracy. By using Eq. (19), the period of oscillation is estimated to be $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 9.4\ {\rm fm}$, where $k_{\rm FP}/\Lambda\sim 0.154$ is the momentum of the Fermi point.
@@ III.4 High-density region
When the chemical potential is large enough, the DCDW phase ($q\neq 0$) is realized. Figure 9 shows the dispersion relation of the quark fields along the momentum $k_{z}$ for fixed parameters $(M/\Lambda,q/\Lambda,\mu/\Lambda)=(0.09,0.62,0.52)$ with $a\Lambda=1$.
Figure 9: Dispersion relations on the lattice with $a\Lambda=1$ in the high-density region ($\mu/\Lambda=0.52$).
In this case, the two touching points of $\omega_{-}$ and $\tilde{\omega}_{-}$ are regarded as the Weyl points. Furthermore, the $\omega_{+}$ and $\omega_{-}$ cross the Fermi level. Thus, there are six nondifferentiable points. In the low-momentum region, the total dispersion relation becomes a flat band due to the cancellations between $\omega_{-}$ and $\tilde{\omega}_{-}$ and between $\omega_{+}$ and $\tilde{\omega}_{+}$, and the edges of the flat band of the total dispersion relation are nondifferentiable. Furthermore, due to the four Fermi points, the total dispersion relation is nondifferentiable at the four momenta.
@@
Figure 10 shows the results for the Casimir coefficient. For the Dirac-sea effect, the Casimir effect shows oscillating behavior, as expected from the Weyl points. Furthermore, we see that the Fermi-sea part of Casimir coefficient oscillates with higher frequency than that from the Dirac sea. The frequencies of the oscillations are determined by the positions of the nondifferentiable points on the dispersion relation. By using Eq. (19), the oscillation period is estimated to be $L_{z}^{\rm osc}=2\pi/k_{\rm WP}\sim 4.8$ fm with $k_{\rm WP}/\Lambda\sim 0.298$ for the Dirac-sea part. For the Fermi-sea part of $\omega_{-}$, the period is determined by $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 1.7$ fm with $k_{\rm FP}/\Lambda\sim 0.851$ in addition to $L_{z}^{\rm osc}=2\pi/k_{\rm WP}\sim 4.8$ fm. For the Fermi-sea part of $\omega_{+}$, $L_{z}^{\rm osc}=2\pi/k_{\rm FP}\sim 7.6$ fm with $k_{\rm FP}/\Lambda\sim 0.190$. In fact, the Dirac-sea and Fermi-sea contributions shown in Fig. 10 are dominated by the periods of $L_{z}^{\rm osc}\sim 4.8$ fm and $L_{z}^{\rm osc}\sim 1.7$ fm, respectively, and the total result is their superposition.